# Stabilization of active tissue deformation by a dynamic morphogen gradient

### v1

@@ VI Discussion
In developing animals, biological tissues can undergo active anisotropic deformation [1]. However, classical active matter results predict that such deformations are unstable under certain conditions [24, 25, 50]. In several systems, the local direction of active tissue deformation is controlled by the gradient of a protein concentration field [12, 30, 7]. Such proteins are called morphogens. They are typically produced by the cells in a localized source region of the tissue, and they can form gradients through diffusion and degrade [1]. However, both morphogen pattern and source region are also advected by the flows generated by the active deformation [37]. Here, we studied the stability of a system which includes source field, morphogen field, and morphogen-gradient-controlled active tissue deformation. Similar to what we found earlier without explicit source region [32], we find that systems where the active stresses are gradient-extensile can be marginally stable, while gradient-contractile systems are always unstable. This is consistent with an observed abundance of gradient-extensile systems in the biological literature [5, 29, 51, 3, 44, 7, 41, 53, 54, 46] and the virtual absence of gradient-contractile systems (also discussed in [55]), suggesting that the instability discussed here may act as an evolutionary selection criterion.
@@
We further find that in the gradient-contractile case, neither degradation nor diffusion can stabilize the system. Degradation only acts – in conjunction with production – to decrease the difference between morphogen and source fields. Meanwhile the system is still subject to the instability. Also, while diffusion can decrease the growth rates of perturbations in the gradient-contractile case, we find that it never suffices to fully stabilize the system. This is because both morphogen and source fields advect with the active flows, while only the morphogen diffuses. In the case of a linear source profile, the maximal perturbation growth rate could become arbitrarily small for sufficiently large diffusion coefficient. However, in the case of a localized source, the maximal perturbation growth rate is always at least $\approx 2$ times the free deformation rate.
We also identify another instability that is not related to the classical active matter instability, which we call “affine diffusion-shear instability”, and which also appears in the gradient-extensile case. More precisely, this instability corresponds to a positive local Lyapunov exponent of a dynamically changing state. It arises from the interplay between the pure shear of the system and the amplitude of the morphogen profile, and in tissues with a localized source whose width is much smaller than the morphogen diffusion length $\sqrt{D/k_{\mathrm{d}}}$. In this case, the extent of the morphogen profile is defined by the diffusion length, while its amplitude is defined by the width of the source region. As a consequence, when the system shear is perturbed to be slightly larger, and consequentially the source width increases, thus the amplitude of the morphogen field increases, which promotes a faster shear. In the context of deforming biological tissues, a positive Lyapunov exponent means that the time-dependent deformation of such tissues is expected to be not very reproducible. In many developing tissues, the morphogen profile has a larger extent than the source region [1, 33, 5], and our work suggests that this instability should occur. This raises the question of how this instability is prevented in developing systems. One possibility might be that there is a saturation in the active stresses as a function of the magnitude of the morphogen gradient. Such a saturation could be mediated for instance through cell polarity [32]. Whether this could be sufficient to suppress this instability is an interesting question for future research.
@@
Our study further motivates experiments, and a key question is whether there are at all any gradient-contractile biological systems. There are many systems where morphogen profiles are known to be necessary for active anisotropic tissue deformation. However, it is most often not yet known whether the gradient of the morphogen controls the directionality or only the magnitude of the active stresses. Thus, experiments are needed that disentangle both. For instance, to modify the direction of the morphogen pattern while keeping the overall magnitude the same, one could combine combine null-mutants with a graded over-expression [30] or the external supply of a gradient, e.g. through morphogen-soaked beads [27, 1]. We believe that combining insights from active matter theory with experiments in this way will help decode the biological mechanisms underlying oriented tissue deformation and their robustness.
@@ A.5 Fourier transform
as in Eq. (37), we have the usual derivation rule, where the Fourier transform of $\mathrm{bar}\partial_{j}f(\bm{\mathrm{bar}r},\mathrm{bar}t)$ is $i\mathrm{bar}{k}_{j}f(\bm{\mathrm{bar}q},\mathrm{bar}t)$. From Eqs. (37) and (28) also follows that a given co-deforming Fourier mode with wave vector $\bm{\mathrm{bar}{q}}$ corresponds to a lab frame Fourier mode with wave vector $\bm{q}$ with components
@@ Appendix B Dynamics in co-deforming coordinates
$\mathrm{bar}\partial_{t}s+\mathrm{bar}{v}_{i}(\mathrm{bar}\partial_{i}s)$ $=0$ (41)
$\mathrm{bar}\partial_{t}c+\mathrm{bar}{v}_{i}(\mathrm{bar}\partial_{i}c)$ $=\big{(}Ds^{-2}_{ij}\mathrm{bar}\partial_{i}\mathrm{bar}\partial_{j}-k_{\mathrm{d}}\big{)}c+k_{\mathrm{d}}s$ (42)
00 0 $=\frac{1}{4}s_{jl}\partial_{i}^{2}\mathrm{bar}{\mathrm{v}}_{l}-\partial_{j}\Pi^{\prime}+\partial_{i}\tilde{\sigma}^{a}_{ij}$ (43)
@@ B.1 Non-affine flows
For deforming systems, we will generally consider the dynamics of source and morphogen fields in co-deforming coordinates, Eqs. (41) and (42). To obtain the non-affine flows, we solve Eqs. (43) and (44) by taking the lab-frame Fourier transform, which yields after some transformations:
@@ B.2 Absence of non-affine flows for $\partial_{y}c=0$
If the morphogen concentration field $c$ is homogeneous in $y$ direction, i.e. $\partial_{y}c=\partial_{\mathrm{bar}{y}}c=0$, then there are no non-affine flows, i.e. there are no flows in co-deforming coordinates, $\bm{\mathrm{bar}{\mathrm{v}}}=0$. For fixed system dimensions, this corresponds to a complete absence of flows, $\bm{{\mathrm{v}}}=0$.
@@ C.2 In co-deforming Fourier space coordinates
[…]
@@ C.3 Numerical solution
To numerically compute the convolutions in Eqs. (53), (54), and (60), we need to take into account that the numerical real-space representations of our fields are spatially discretized. To discuss the main ideas, we will discuss here the case of a single spatial dimension, $\mathrm{bar}{x}$, in co-deforming coordinates. The generalization to lab frame coordinates and/or several dimensions will be straight forward. Specifically, we consider the periodic interval $\mathrm{bar}{x}\in[-L_{x}(0)/2,+L_{x}(0)/2)$ to be discretized by $N_{x}$ equal-length steps $\Delta\mathrm{bar}x=L_{x}(0)/N_{x}$, where $N_{x}$ is an even integer. Any function $f(\mathrm{bar}{x})$ is represented by a set of numbers $f_{k}$ with $k=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$, such that $f_{k}=f(k\Delta\mathrm{bar}x)$. As a consequence, the Fourier series are not infinite any more as in Eq. (37), but instead cut off at some finite wave vectors. In particular, the numerical co-deforming Fourier transform of $f$ is represented by a set of complex numbers $\tilde{f}_{j}$ with $j=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$, where $\tilde{f}_{j}=f(\mathrm{bar}{q}_{x}=j\mathrm{bar}{q}_{0})$ with $\mathrm{bar}{q}_{0}:=2\pi/L_{x}(0)$. Eq. (37) then becomes:
@@
where we used that $\mathrm{bar}{q}_{0}\Delta\mathrm{bar}x=2\pi/N_{x}$. To maintain in this discretized space the property that the Fourier transform of the product of two functions $f$ and $g$ is the convolution of their respective Fourier transforms,
$\widetilde{(fg)}_{j}=(\tilde{f}\ast\tilde{g})_{j},$ (62)
we need to define the following discrete convolution operator:
@@
for $j=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$. Yet, we see that the index $j-j^{\prime}$ of $\tilde{g}$ can get outside of the range $-N_{x}/2,-N_{x}/2+1,\dots,N_{x}/2-1$. For Eq. (62) to still be fulfilled, we need to define $\tilde{g}$ to be periodic, such that $\tilde{g}_{(j+nN_{x})}:=\tilde{g}_{j}$ for all $n\in\mathbb{Z}$. In this sense, the convolution in Fourier space of our discretized function is circular. We use such a circular definition to numerically compute the convolutions in Eqs. (53), (54), and (60), to preserve the relation Eq. (62).
@@ Appendix D Fixed-size system with linear source
For a linear source and morphogen fields, the active flows, Eq. (60), simplify to:
[…]
@@ D.1 No diffusion, $D=0$
$\delta b$ $=\delta c+\frac{k_{\mathrm{d}}}{\omega_{\mathrm{act}}}\delta s.$ (72)
@@
$\partial_{t}\delta b$ $=\omega_{\mathrm{act}}\partial_{t}\delta b.$ (74)
Hence, the difference between both morphogen and source, $\delta\Delta$, decays at rate $k_{\mathrm{d}}$, while a superposition of both, $\delta b$, grows with the same growth rate as the system without source [32].
@@ E.1 Value of $s_{b}$ in dimensionless units
Figure 8: Linear stability of gradient-extensile systems with $\mathrm{{w}}\rightarrow\infty$, i.e., where the source profile converges to a single cosine mode. Green dots and green circles represent the numerical analysis with and without discrete periodic convolution, respectively. Inset showing that perturbation modes with $\omega>0$ disappear when convolution is not periodic.
E.2 Numerically growing modes in gradient-extensile systems
In section IV, we found that in a gradient-extensile systems, some modes showed a positive growth rate. Yet, when we increased the spatial discretization $N_{x}$, the value of the growth rate decreased as a power law with $N_{x}$ (Figure 3H), suggesting that these positive growth rates may just be a numerical artifact. Here we discuss this question further.
We discuss the positive growth rates for the limit $\mathrm{{w}}\rightarrow\infty$, for which the source profile $s_{0}(\bm{r})$, defined in Eq. (20), becomes a cosine. We obtained the modes with positive growth rates, using the approach described in Appendix C.3. Yet, for this we used the circular convolutions, Eq. (63), which in a way is a natural choice, imposed by the fact that we do not have infinitely fine resolution of $s(x)$ and $c(x)$, which would allow to have the information about these fields at arbitrarily large wave number $\bm{q}$. Instead, because these fields are discretized, the Fourier amplitude at high wave vectors just become periodic, as described in Appendix C.3. In other words, due to a limited resolution, we do not know what the amplitude for high wave vectors is and we “just” set them to some values assuming periodicity. One could also use a different approach, by assuming that the Fourier amplitudes at these high wave vectors just vanishes. This corresponds to an alternative numerical implementation of the Fourier-space convolution. Specifically, we replace the circular convolutions by “non-circular” convolutions where we ignore the periodicity of the discrete Fourier transforms, i.e., we replace Eq. (63) by:
$(\tilde{f}\ast\tilde{g})_{j}=\sum_{j^{\prime}=-N_{x}/2}^{N_{x}/2-1}{\tilde{f}_{j^{\prime}}\tilde{g}_{j-j^{\prime}}\Theta_{j-j^{\prime}}}$ (79)
where
$\Theta_{k}=\begin{cases}1&\text{for $-N_{x}/2\leq k\leq N_{x}/2-1$,}\\ 0&\text{otherwise.}\end{cases}$ (80)
The factor $\Theta_{j-j^{\prime}}$ explicitly ensures that there is any periodicity in $\tilde{g}_{j-j^{\prime}}$ is ignored. We compare the solutions of the eigenvalue problem, Eq. (67), between circular and non-circular convolution in Figure 8, and we find that while most of the modes did not change much, the modes with positive growth rates disappeared when using the non-circular convolutions. This further suggests that these positive growth rates are merely an artifact of our spatial discretization.
Figure 9: System dynamics of a freely deforming system with a diffusive morphogen, $\tilde{v}_{xx}^{0}=\tilde{v}_{xx}^{\mathrm{free}},D=10^{-2}$. (A) Morphogen magnitude over time for the gradient-extensile case, where we compare the dynamic system (green) with the stationary value (Eq. (83), dashed red) as a function of time (B) Absolute values of shear rates of the systems in panel A. (C,D) Morphogen and shear rate magnitudes for the gradient-contractile case. (E) Decomposition of the partial time derivative of the dynamic morphogen amplitude in the gradient-contractile case (see text).

### v2

@@ VI Discussion
In developing animals, biological tissues can undergo active anisotropic deformation [1]. However, classical active matter results predict that such deformations are unstable under certain conditions [24, 25, 53]. In several systems, the local direction of active tissue deformation is controlled by the gradient of a signaling molecule concentration field [12, 33, 7]. Such signaling molecules are typically produced by the cells in a localized source region of the tissue, and they can form gradients through diffusion and degrade [1]. However, both signaling pattern and source region are also advected by the flows generated by the active deformation [36]. Here, we study the stability of a system which includes source field, signaling field, and signaling-gradient-controlled active tissue deformation. We find that systems where the active stresses are gradient-contractile are always unstable, while gradient-extensile can be marginally stable under certain conditions. Our findings are consistent with, and potentially explain, an observed abundance of gradient-extensile systems in the biological literature [5, 32, 54, 3, 47, 7, 44, 56, 57, 49] and the virtual absence of gradient-contractile systems (also discussed in [58]), suggesting that the instability discussed here may act as an evolutionary selection criterion.
@@
In our previous work, we studied a simplified signaling dynamics, where the overall gradient was imposed by the boundary conditions [35]. Here, we show that more realistic signaling dynamics, including secretion, diffusion, and degradation, lead to a number of important differences. First, in the gradient-contractile case, neither degradation nor diffusion can stabilize the system. While diffusion can decrease the growth rates of perturbations, we find that it never suffices to fully stabilize the system. This is because both signal and source fields are advected with the active flows, while only the signal diffuses. In the case of a linear source profile, the maximal perturbation growth rate could become arbitrarily small for a sufficiently large diffusion coefficient. However, in the case of a localized source, the maximal perturbation growth rate converges to a finite constant in the limit of a large diffusion coefficient (Appendix E.2).
Second, we show that for a localized source region, gradient-extensile systems can become unstable whenever the tissue is too short in the direction of the gradient as compared to the transversal direction. Conversely, we numerically obtain marginal stability whenever the tissue is more elongated in the direction of the gradient or when the signaling lengths scale, $\sqrt{D/k_{\mathrm{d}}}$, is small. However, we also find that even in the unstable regime, the maximum growth rate of gradient-extensile systems are often orders of magnitude smaller than the free deformation rate, in particular for aspect ratios close to one (Figs. 3E-J). It is not impossible that during development, biological tissues transiently accept an instability with a small growth rate until entering a stable regime once they become more elongated.
Finally, in deforming systems, an “affine diffusion-shear instability” can appear. Different from the other instabilities discussed here, which are all related to the Simha-Ramaswamy instability, the affine diffusion-shear instability does not require non-affine tissue flows. Instead, it is only reflected in an interplay between the amplitude of the signaling molecule and the overall tissue shear rate. This instability appears in tissues with a localized source whose width is smaller than the signaling diffusion length $\sqrt{D/k_{\mathrm{d}}}$. It occurs even when adding any non-linearity in the way active stress depends on the signaling gradient (Appendix G). In many developing tissues, the signaling profile has a larger extent than the source region [1, 38, 5], and our work suggests that this instability should occur. Although this instability does not perturb the patterns of source, signaling molecule, or flow, it perturbs their amplitude and, most importantly, leads to a lack of control of the final tissue dimensions. This points to an additional feedback mechanism not included in our model that may be required to control the final tissue shape.
@@
Our study further motivates experiments, and a key question is whether there are at all any gradient-contractile biological systems. There are many systems where signaling gradients are known to be necessary for active anisotropic tissue deformation. However, in most systems, it is not yet known whether the signal controls the directionality or only the magnitude of the active stresses. For instance, there are still open debates about what controls the directionality of active stresses in what is probably the best-studied anisotropic tissue deformation process, which is germ band extension (GBE) in the fruit fly Drosophila melanogaster. Recent work on GBE proposed that a dorso-ventral signaling gradient modulates the magnitude of junctional myosin, which plays the role of active stresses [59]. Yet, different mechanisms have been proposed for the control of the myosin anisotropy and its directionality. Classically, these include anterior-posterior-oriented gradients [44, 13], which would correspond to a gradient-extensile coupling. But more recently, a group has proposed that a dorso-ventrally aligned tissue tension could provide a global mechanical orienting signal [36, 59]. Thus, in many cases more experiments are needed to cleanly disentangle the signals that merely modulate the amplitude of active stresses from those that control their direction. For instance, to modify the direction of the signaling pattern while keeping a similar overall magnitude, one could combine null-mutants with a graded over-expression [33] or the external supply of a gradient, e.g. through signaling-molecule-soaked beads [27, 1]. We believe that systematically combining insights from active matter theory with experiments in this way will help decode the biological mechanisms underlying oriented tissue deformation and its robustness.
@@ A.5 Fourier transform
As in Eq. (35), we have the usual derivation rule, where the Fourier transform of $\mathrm{bar}\partial_{j}f(\bm{\mathrm{bar}r},\mathrm{bar}t)$ is $i\mathrm{bar}{k}_{j}f(\bm{\mathrm{bar}q},\mathrm{bar}t)$. From Eqs. (35) and (26), it also follows that a given co-deforming Fourier mode with wave vector $\bm{\mathrm{bar}{q}}$ corresponds to a lab-frame Fourier mode with wave vector $\bm{q}$ with components
@@ Appendix B Dynamics in co-deforming coordinates
$\mathrm{bar}\partial_{t}s+\mathrm{bar}{{\mathrm{v}}}_{i}(\mathrm{bar}\partial_{i}s)$ $=0$ (39)
$\mathrm{bar}\partial_{t}c+\mathrm{bar}{{\mathrm{v}}}_{i}(\mathrm{bar}\partial_{i}c)$ $=\big{(}Ds^{-2}_{ij}\mathrm{bar}\partial_{i}\mathrm{bar}\partial_{j}-k_{\mathrm{d}}\big{)}c+k_{\mathrm{d}}s$ (40)
$0$ $=\frac{1}{4}s_{jl}\partial_{i}^{2}\mathrm{bar}{\mathrm{v}}_{l}-\partial_{j}\Pi+\partial_{i}\tilde{\sigma}^{a}_{ij}$ (41)
@@ B.1 Non-affine flows
For deforming systems, we will generally consider the dynamics of source and signaling fields in co-deforming coordinates, Eqs. (39) and (40). To obtain the non-affine flows, we solve Eqs. (41) and (42) by taking the lab-frame Fourier transform, which yields after some transformations:
@@ B.2 Absence of non-affine flows for $\partial_{y}c=0$
If the signaling field $c$ is homogeneous in $y$ direction, i.e. $\partial_{y}c=\partial_{\mathrm{bar}{y}}c=0$, then there are no non-affine flows, i.e. there are no flows in co-deforming coordinates, $\bm{\mathrm{bar}{\mathrm{v}}}=0$. For fixed system dimensions, this corresponds to a complete absence of flows, $\bm{{\mathrm{v}}}=0$.
@@ C.2 In co-deforming Fourier space coordinates
$\delta\tilde{\sigma}^{a}_{xj}=\frac{i\mathrm{sgn}(\alpha)}{l_{x0}}c_{0}^{\prime}\ast(q_{j}\delta c)\quad\text{for {\hbox{j\in\{x, y\}}}}.$ (56)
@@ C.3 Numerical solution
To numerically compute the convolutions in Eqs. (51), (52), and (58), we need to take into account that the numerical real-space representations of our fields are spatially discretised. To discuss the main ideas, we will discuss here the case of a single spatial dimension, $\mathrm{bar}{x}$, in co-deforming coordinates. The generalisation to lab frame coordinates and/or several dimensions will be straightforward. Specifically, we consider the periodic interval $\mathrm{bar}{x}\in[-L_{x}(0)/2,+L_{x}(0)/2)$ to be discretised by $N_{x}$ equal-length steps $\Delta\mathrm{bar}x=L_{x}(0)/N_{x}$, where $N_{x}$ is an even integer. Any function $f(\mathrm{bar}{x})$ is represented by a set of numbers $f_{k}$ with $k=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$, such that $f_{k}=f(k\Delta\mathrm{bar}x)$. As a consequence, the Fourier series are not infinite any more as in Eq. (35), but instead cut off at some finite wave vectors. In particular, the numerical co-deforming Fourier transform of $f$ is represented by a set of complex numbers $\tilde{f}_{j}$ with $j=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$, where $\tilde{f}_{j}=f(\mathrm{bar}{q}_{x}=j\mathrm{bar}{q}_{0})$ with $\mathrm{bar}{q}_{0}:=2\pi/L_{x}(0)$. Eq. (35) then becomes:
@@
where we used that $\mathrm{bar}{q}_{0}\Delta\mathrm{bar}x=2\pi/N_{x}$.
We define the convolution of two Fourier transforms, $\tilde{f}_{j}$ and $\tilde{g}_{j}$ as follows:
@@
for $j=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$. Yet, we see that the index $j-j^{\prime}$ of $\tilde{g}$ can get outside of the range $N_{x}/2,-N_{x}/2+1,\dots,N_{x}/2-1$. To prevent aliasing effects, we set $\tilde{g}_{j}\equiv 0$ whenever $j<-N_{x}/2$ or $j\geq N_{x}/2$. Note that this technique for de-aliasing is akin to the zero-padding method [60, 61], frequently used for spectral solvers.
@@ Appendix D Fixed-size system with linear source
For linear source and signaling fields, the active flows, Eq. (58), simplify to:
$\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c\qquad$ (65)
with
$\omega_{\mathrm{act}}=4\mathrm{sgn}(\alpha)\sin^{2}{\phi},$ (66)
@@ D.1 No diffusion, $D=0$
$\delta c_{\mathrm{eff}}$ $=\delta c+\frac{k_{\mathrm{d}}}{\omega_{\mathrm{act}}}\delta s.$ (70)
@@
$\partial_{t}\delta c_{\mathrm{eff}}$ $=\omega_{\mathrm{act}}\delta c_{\mathrm{eff}}.$ (72)
Hence, the difference between both signaling and source, $\delta\Delta$, decays at rate $k_{\mathrm{d}}$, while a superposition of both, $\delta c_{\mathrm{eff}}$, grows with the same growth rate as the system without source [35].
@@ Appendix E Fixed-size system with localized source
Figure 5: Instability in a gradient-extensile system with fixed size and a localized source profile. (A) Snapshots of the late-time dynamics of the system in Figure 3C, showing the source profile at $t=300$ (left) and $t=1500$ (right). (B) Perturbation mode with maximum growth rate for system from Figure 3C and panel A. The color plot shows the signaling molecule perturbation, $\delta c$, and white arrows show the flow field perturbation, $\delta\bm{{\mathrm{v}}}$.
@@ E.1 Value of $s_{b}$ in dimensionless units
E.2 Convergence to a finite perturbation growth rate for $D\rightarrow\infty$
Figure 6: Largest perturbation growth rates $\omega_{\mathrm{max}}$ of a fixed-size system as a function of the signaling molecule diffusion. Parameter values, $k_{\mathrm{d}}=1$ and $L_{x}=L_{y}=1$.
The large-diffusion limit can be understood similarly to section III, where we assume that $Dq_{\mathrm{min}}^{2}$ is faster than any other rate to eliminate the signaling perturbation dynamics. Specifically, only the source perturbation dynamics, Eq. (51), remains, where the signaling perturbation $\delta c$ depends directly on the source perturbation through Eq. (19), which implies $\delta c\sim D^{-1}\delta s$. Moreover, for large $D$, Eq. (22) implies that $c_{0}\sim D^{-1}s_{0}$, and thus $c_{0}^{\prime}\sim D^{-1}s_{0}^{\prime}$.
In dimensionless units, we set $s_{b}$ such that the free deformation rate is one, which implies $s_{0}\sim s_{b}\sim D$ (compare Eq. (75)). Thus, we have $c_{0}^{\prime}\sim D^{0}$. As a consequence, the perturbation of active stresses and active flows scale as $\delta\mathrm{bar}{\mathrm{v}}_{x}\sim\delta\tilde{\sigma}^{a}_{ij}\sim c_{0}^{\prime}\delta c\sim D^{-1}$ (compare Eqs. (56) and (58)). The active flows enter in the advective term of the source field, Eq. (51), where they are multiplied with the stationary source field, which scales as $s_{0}^{\prime}\sim s_{b}\sim D$. Hence, in the large-$D$ limit, the source dynamics does not depend on $D$ any more. As a consequence, the largest perturbation growth rate is constant.
In dimensionful units, i.e. if we imposed a constant source magnitude $s_{b}\sim D^{0}$, we would have $c_{0}^{\prime}\sim D^{-1}$. Hence, the perturbation growth rate would scale as $c_{0}^{\prime}\delta c\sim D^{-2}$, and the free deformation rate would scale as $\sim\tilde{\sigma}^{a}_{xx}\sim(c_{0}^{\prime})^{2}\sim D^{-2}$. Hence, the perturbation growth rate decreases with $D$ in the same way as the free deformation rate.